Light is incident from air onto the surface of a liquid. The angle of incidence is 56.2
Answer #1

from snells law

n1sin i = n2 sin r

a sin56.2 = n2 sin35.4

n2 =0.697

tanq = n1/n2 = 1/0.697

q = 55.11 degrees